### v1

# PathDB: A system for evaluating regular path queries

Abstract.
PathDB is a Java-based graph database designed for in-memory data loading and querying. By utilizing Regular Path Queries (RPQ) and a closed path algebra, PathDB processes paths through its three main components: the parser, the logical plan, and the physical plan. This modular design allows for targeted optimizations and modifications without impacting overall functionality. Benchmark experiments illustrate PathDB’s execution times and flexibility in handling dynamic and complex path queries, compared to baseline methods like Depth-First Search (DFS) and Breadth-First Search (BFS) guided by an automaton, highlighting its optimizations that contribute to its performance.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing
Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/dbgutalca/PathDB.
1. Introduction
In the rapidly evolving field of graph databases, several systems have emerged, each showcasing distinct strengths and capabilities. Among the most prominent players are Neo4j, Kuzu and Memgraph. While these systems support path querying, they share a common characteristic: the reliance on specific algorithms to process recursive paths queries. However, Neo4j, Kuzu, and Memgraph offer only limited support for path queries.
PathDB focuses on utilizing execution trees instead of recursive algorithms, as used by most systems. This approach allows for future optimization of these trees through techniques such as predicate pushdown and other enhancements. Its uniqueness lies in a path manipulation algebra, inspired by relational algebra, based on the concept of sets of paths. This path algebra consists of three main operators: selection, which filters paths; join, which concatenates compatible paths; and union, which merges sets of paths. It is worth noting that all these operators function exclusively on sets of paths.
PathDB’s design features three main components: the parser, the logical plan, and the physical plan. By utilizing a path query pattern inspired by the GQL standard (2024iso, ), PathDB enables users to query paths and return a set of paths as results.
The primary objective of this paper is to showcase the innovative features and capabilities of PathDB. We provide a comprehensive overview of how graph data is stored, the query evaluation workflow, the syntax and operation of its query language, and related optimizations. Additionally, we demonstrate how PathDB’s design allows for targeted optimizations and modifications without disrupting overall functionality. This includes optimizations in the logical plan, changes to the physical plan, and the ability to seamlessly transition between different storage systems.
Section 2 provides a system overview, describing the architecture and data storage. Section 3 presents a demonstration of PathDB’s interface and usage, including a comparison with base methods like DFS and BFS with an automaton. Finally, Section 4 offers conclusions and future work.
Figure 1. A graph representing a social network (drawn from the LDBC SNB benchmark) (EDBT.2025.63, ).
2. System overview
Figure 2. Query evaluation workflow.
{forest}
for tree=scale=0.9, align=center [$\pi_{y.name}^{10}$ [$\sigma_{first.name=\texttt{"Moe"}}$ [$\cup$ [$\phi^{{\small\texttt{isTrail}}}$ [$\sigma_{label(edge(1))=\texttt{Knows}}$ [$\texttt{Paths}_{1}(G)$, yshift=-1.1cm] ] ] [$\phi^{{\small\texttt{isTrail}}}$ [$\bowtie^{{\small\texttt{isTrail}}}$ [$\sigma_{label(edge(1))=\texttt{Likes}}$ [$\texttt{Paths}_{1}(G)$] ] [$\sigma_{label(edge(1))=\texttt{HasCreator}}$ [$\texttt{Paths}_{1}(G)$] ] ] ] ] ] ]
(a) Basic query plan.
{forest}
for tree=scale=0.9 [$\pi_{y.name}^{10}$ [$\cup$ [$\phi^{{\small\texttt{isTrail}}}$ [$\sigma_{first.name=\texttt{"Moe"}}$ [$\sigma_{label(edge(1))=\texttt{Knows}}$ [$\texttt{Paths}_{1}(G)$, yshift=-1cm] ] ] ] [$\phi^{{\small\texttt{isTrail}}}$ [$\bowtie^{{\small\texttt{isTrail}}}$ [$\sigma_{first.name=\texttt{"Moe"}}$ [$\sigma_{label(edge(1))=\texttt{Likes}}$ [$\texttt{Paths}_{1}(G)$] ] ] [$\sigma_{label(edge(1))=\texttt{HasCreator}}$ [$\texttt{Paths}_{1}(G)$] ] ] ] ] ]
(b) Optimized query plan.
Figure 3. A query plan for the query “MATCH TRAIL p = (x)-[(Knows|Likes.HasCreator)+]->(y) WHERE x.name = "Moe" Return y.name Limit 10”.
In this section, we provide a comprehensive overview of the principal PathDB features. We will cover how the data is stored, the architecture’s design, the operation of its query language, and related optimizations.
It is worth noting that PathDB is a Java-based system designed to load and query data in memory. It features an architecture with three main components: the parser, the logical plan, and the physical plan. By using a path query pattern based on Regular Path Queries (RPQ), PathDB enables users to query for paths and return a set of paths as results. Importantly, PathDB employs a closed path algebra, meaning that each operator takes a set of paths as input and returns a set of paths as output.
Graph Storage
As previously mentioned, PathDB is a system that maintains graph data in memory. Consequently, it is necessary to load this data into memory with each use.
The information about nodes and edges is stored in a data structure that uses a compressed sparse row with vertical partitioning (CSR VP) as its base. Essentially, a CSR is represented by two lists: the first list maintains the offsets into the second array, marking the beginning of the outgoing edges for the corresponding vertex. The end of this list is marked by the beginning of the next vertex. The second array contains the edge identifiers (Tetzel2018, ). Combining CSR with vertical partitioning (VP) allows us to leverage the characteristics of a CSR while using labels, which is beneficial for our case since Regular Path Queries make use of these labels (Tetzel2018, ).
In our implementation, we have two hashmaps: one where the key is the label and the value is a Node, and another where the key is the label and the value is a linked list of Edges that contain this label. It is worth mentioning that, so far, we have demonstrated how the graph is stored, but not how to work with paths. Since the logical and physical operators in PathDB handle sets of paths, it is necessary to transform the nodes and edges of a graph $G$ into sets of paths. To achieve this, two main functions are used: $\texttt{Paths}_{0}(G)$, which converts the set of nodes into paths of length 0, and $\texttt{Paths}_{1}(G,l)$, which transforms the edges of a graph $G$ with a specific label $l$ into a set of paths of length 1.
Path Algebra
Our path algebra (EDBT.2025.63, ) consists of four fundamental operators, which can be composed and evaluated over graph-structured data. The selection operator ($\sigma$) filters paths by applying conditions over nodes or edges. The union operator ($\cup$) merges two sets of paths, $a$ and $b$, into a single set. The path join operator ($\bowtie^{\tau}$) concatenates paths from set $a$ with compatible paths from set $b$; the join may optionally be constrained by a path semantic $\tau$, which can be one of isTrail for no repeated edges, isAcyclic for no repeated nodes, or isSimple for no repeated nodes, except the first and last node. Finally, the recursive operator ($\phi^{\tau}$) applies repeated join operations to compute transitive expansions of a path set, either until a fix point is reached or until specific criteria are satisfied. Like the join, the recursive operator can also be constrained by a path semantic $\tau$, allowing it to generate only trails, simple, or acyclic paths when required. These operators form the foundation for expressing rich navigational queries. An example that demonstrates the combined application of selection, union, join, and recursion is presented in Figure 1, shaped as an execution tree.
Path Query Pattern
A path query pattern is an expression inspired by the GQL standard (2024iso, ), and consists of three main components: the MATCH clause, that defines the path pattern, the WHERE clause, that defines the selection conditions, and the RETURN clause, that specifies the output.
        MATCH <restrictor>? <var> = <pathPattern>
        (WHERE <conditions>)?
        RETURN <term>+ <limit>?
The MATCH clause consists of three components: a path restrictor (WALK, TRAIL, ACYCLIC, or SIMPLE), a path variable var (a symbolic identifier formed by the concatenation of characters), and a path pattern pathPattern. The path pattern has the form (x)-[r]->(y), where x and y represent the source and target variables (these can be explicitly named and later referenced in subsequent clauses), and r is a regular expression over the edge labels of the graph. Let $r_{1}$ and $r_{2}$ be regular expressions. Valid constructions for r include: a label $r_{1}$; a negated label $!r_{1}$; concatenation $r_{1}r_{2}$; alternation $r_{1}\mid r_{2}$; Kleene star $r_{1}^{*}$; positive closure $r_{1}^{+}$; and the optional expression $r_{1}?$.
The WHERE clause consists of a set of selection condition terms used to filter the paths retrieved in the MATCH clause. Let $i\geq 1$ be a natural number, pr a property label, $v$ a value, $\oplus$ a comparison operator ($=$, $\neq$, $<$, $>$, $\leq$, $\geq$), and sVar and tVar symbolic variables representing the source and target nodes of a path pattern. A selection condition $c$ is defined recursively. A simple selection condition in PathDB can be any expression such as sVar.pr $\oplus$ v, tVar.pr $\oplus$ v, FIRST().pr $\oplus$ v, LAST().pr $\oplus$ v, NODE(i).pr $\oplus$ v, EDGE(i).pr $\oplus$ v, LABEL(sVar) $\oplus$ v, LABEL(tVar) $\oplus$ v, LABEL(FIRST()) $\oplus$ v, LABEL(LAST()) $\oplus$ v, LABEL(NODE(i)) $\oplus$ v, LABEL(EDGE(i)) $\oplus$ v, LENGTH() $\oplus$ v, as well as boolean predicates such as ISTRAIL(), ISSIMPLE(), and ISACYCLIC(). If $c_{1}$ and $c_{2}$ are selection conditions, then $(c_{1}{\small\texttt{ AND }}c_{2})$ and $(c_{1}{\small\texttt{ OR }}c_{2})$ are complex selection conditions.
Finally, the RETURN clause specifies the elements of the path to be projected in the result. Projection expressions may reference any variable defined in the MATCH clause, including path and node variables. Let $i\geq 1$ be a natural number, pr a property label, Var a path pattern variable, and sVar and tVar the symbolic identifiers of the source and target nodes of that path. A PathDB projection term can refer to node or edge attributes using expressions such as sVar.pr, tVar.pr, NODE(i).pr, EDGE(i).pr, FIRST().pr, and LAST().pr. It can also include functions that extract structural or semantic features of the path, such as LABEL(), LENGTH(), ISTRAIL(), ISSIMPLE(), and ISACYCLIC(). An optional LIMIT clause may be appended to restrict the number of returned paths.
System Design
The PathDB design is structured around three primary components, reflecting its query evaluation workflow, as shown in Figure 2: the parser, the logical plan, and the physical plan.
First, the parser is responsible for validation, extraction, and transformation of a path query pattern into an abstract syntax tree (AST). This process is carried out using ANTLR4111ANTLR (ANother Tool for Language Recognition), https://www.antlr.org , a tool that, given a specific grammar (in this case, the grammar of a path pattern query for PathDB), converts the query into an AST. This AST has three main components: the source node, the regular expression, and the target node.
Second, the AST, specifically the regular expression, serves as the input for the logical planner component. The AST is then transformed into a logical plan tree. Essentially, we convert the regular expression and endpoint filters (represented as a tree) into multiple algebraic operators. It is worth noting that our logical plan tree is executed in a bottom-up manner, from the leaves to the root.
For example, to retrieve paths that start at a node named $Moe$ and traverse one or more repetitions of either a Knows or Likes edge followed by a HasCreator edge, a user can formulate the query:
MATCH TRAIL p = (x)-[(Knows|Likes.HasCreator)+]->(y) WHERE x.name = "Moe" RETURN y.name LIMIT 10.
The above query is based on the graph of Figure 1, the corresponding logical plan is presented in Figure 3(a). In this logical plan, the tree initially retrieves all edges from the graph $G$ labeled with Knows, Likes and HasCreator, subsequently transforming these edges into a set of paths. If we evaluate the expression from the left-hand side of the union, a recursive operator is applied over the set of paths labeled with Likes. On the right-hand side, a recursive operator is applied over the Join of two path sets: one labeled with Likes and the other with HasCreator. It is important to note that the path semantic isTrail is enforced on the resulting paths—both for the recursion and for the join—ensuring that no edge is repeated during evaluation.
Once the recursion completes, we filter the resulting paths to retain only those whose source node has the property "Name = Moe". Finally, we return the names of the individuals in the last node of each path, limiting the output to 10 results.
As can be observed, since we are working with a tree structure where each operation processes a set of paths, various optimizations can be applied during execution, which could improve execution times. These optimizations are discussed further later on.
Finally, the output of the logical plan is used to retrieve results from the graph through the physical plan. To achieve this, each logical operator in the logical plan is translated into a corresponding physical operator. In this case, the physical plan operates on the graph stored in memory, processing each result on demand.
The design of PathDB allows for the isolation of these three components, thereby enabling optimizations or modifications to be made to each one without causing significant issues. This isolation, for instance, allows for optimizations in the logical plan, changes to the physical plan, such as moving from memory storage to disk storage or any other system, like a database, among other possible modifications.
Current optimizations
Leveraging the Path Algebra to represent logical plans for evaluating Regular Path Query Semantics (RPQS) offers considerable potential for optimization through query rewriting. The algebraic framework enables the systematic transformation of queries into more efficient forms while preserving their semantic integrity. A prominent optimization technique within this context is predicate pushdown, which involves moving constraints on the source or target nodes of paths closer to the leaf nodes in the logical plan. For instance, for the query discussed before, MATCH TRAIL p = (x)-[(Knows|Likes.HasCreator)+]->(y) WHERE x.name = "Moe" RETURN y.name LIMIT 10, Figure 3(a) illustrates the original query plan. Meanwhile, Figure 3(b) presents the optimized plan, where the selection of the source node has been pushed down through the union for each child on the left-hand side.
3. Demonstration
Table 1. - Path Queries and their translation to PathDB over the graph in Figure 1.
Q Query Type RPQ PathDB Query
Q1 People that Bart knows Adjacency Knows MATCH WALK p = (x)-[Knows]->(y) WHERE x.name = "Bart" RETURN y.name
Q2 People who like message Msg1 Adjacency Likes MATCH WALK p = (x)-[Likes]->(y) WHERE y.txt = "Msg1" RETURN x.name
Q3 Who connects Moe to Apu? Reachability Knows+ MATCH TRAIL p = (x)-[Knows+]->(y) WHERE x.name = "Moe" AND y.name = "Apu" RETURN p
Q4 Which messages link Moe to Apu? Reachability (Likes.HasCreator)+ MATCH TRAIL p = (x)-[(Likes.HasCreator)+]->(y) WHERE x.name = "Moe" AND y.name = "Apu" RETURN p
In this section, we demonstrate PathDB’s interface and usage. We showcase how to perform path queries, load data into memory, and interact with the system’s components. Additionally, we provide a comparison with base methods such as Depth-First Search (DFS) and Breadth-First Search (BFS) with an automaton to illustrate PathDB’s capabilities. PathDB is available on GitHub [^2] https://github.com/dbgutalca/PathDB.
Dataset
The dataset is a property graph depicted in Figure 1, which represents a fragment of the graph supplied by the LDBC SNB (SzarnyasWSSBWZB22, ). This graph comprises two types of nodes, labeled as Person and Message, along with three types of edges, labeled as Knows, Likes, and HasCreator. A notable feature of this graph is its ability to support recursive operations, attributed to the presence of cycles. Specifically, the graph includes an inner cycle formed by Knows edges and an outer cycle that traverses the concatenation of Likes and HasCreator edges.
Using PathDB
To load data into PathDB, two files are required: one for nodes and another for edges. Each file follows a structure inspired by the Property Graph Data Format (PGDF) (10734106, ). PathDB introduces two exceptions: the mandatory attribute @dir is always set to T, as PathDB exclusively handles directed edges, and only nodes can have properties. Once the node and edge files are prepared, PathDB can be executed with the command: java -jar PathDB.jar -n nodesFile -e edgesFile, where the node file is always listed first. If the files are supplied in the incorrect order, the default graph will be loaded, corresponding to the graph shown in Figure 1. Maintaining the correct argument order is essential for proper loading and execution.
PathDB provides several configuration options, such as result limits, maximum recursion depth, maximum path length, path semantics, and toggling optimizations on or off. Each configuration setting has an associated command accessible via the help menu (/h). To execute a query in PathDB, the structure specified in the Path Query Pattern from Section 2 must be followed. For instance, an example query could be: MATCH TRAIL p = (x)-[Knows+]->(y) WHERE x.name = "Lisa" RETURN p;. PathDB will then calculate all paths that satisfy the given RPQ (Regular Path Query) according to the defined conditions.
Query results are displayed in the console, with the number of results influenced by PathDB’s configurations. Each result includes the path number and the sequence of node-edge objects that constitute the path. For instance, for the aforementioned query, a possible result could be: Path #3 - p3 e3(Knows) p2 e4(Knows) p4.
Table 1 presents several queries for the graph shown in Figure 1. Each query is defined in textual form, along with its type, the RPQ, and its equivalent PathDB query. The table includes adjacency queries such as Q1 and Q2, which aim to find paths of length 1. On the other hand, it also includes reachability queries, such as Q3 and Q4, which seek paths of undefined length.
4. Experimental evaluation
To demonstrate the functionality of PathDB, we first compare it against a baseline approach based on traditional graph traversal algorithms, specifically, Breadth-First Search (BFS) and Depth-First Search (DFS) with an automaton. We then extend the comparison to include both commercial and academic graph database systems.
Experimental setup
All queries were executed on an Ubuntu Server equipped with 32 GB of RAM. Each query was run three times over the SF1 dataset, with the number of returned results limited to 100 entries and a timeout threshold of 120 seconds per execution.
Dataset
To demonstrate the functionality of PathDB, we used data derived from the LDBC Social Network Benchmark (SNB) (SzarnyasWSSBWZB22, ). Specifically, we selected the dataset at Scale Factor 1 (SF1), which contains 3,181,724 nodes and 17,256,038 edges. This dataset models a realistic social network with rich structural and semantic diversity, enabling a wide range of path-oriented queries. Notably, the SNB graph consists of a static component—unchanged across scale factors—and a dynamic component, which evolves as the graph grows. In our evaluation, we prioritized using labels predominantly present in the dynamic part of the graph to better observe how these variations impact query performance.
4.1. Comparison with baseline
To evaluate PathDB against the baseline algorithms, BFS and DFS with automaton, we selected a set of recursive queries based on the LDBC SNB schema. The queries are presented in the format (x,er,y), where x is the starting node, y is the target node, and er represents the regular expression defining their relationship. For each query, only the starting node was fixed. These queries were: Q1 (p84, (knows+).likes, y), Q2 (f36, hasMember.(knows+), y), Q3 (p84, knows | (knows+), y), Q4 (p3378, (likes.hasCreator)+, y), Q5 (p84, knows+, y), Q6 (p84, (knows+) | likes, y), and Q7 (p10, (workAt | knows)+, y).
Figure 4 compares the execution times of DFS + Automaton, BFS + Automaton, and PathDB. PathDB demonstrates a better performance, completing all queries within 4 seconds, including the complex Q4 , which it processes in 3.945 seconds while DFS and BFS time out at 120 seconds (indicated with an “x”). For simpler queries like Q5, PathDB achieves an execution time of 0.025 seconds, significantly outperforming DFS and BFS, which take over 69 seconds. These results, shown in detail in the Table 2, highlight PathDB’s efficiency and scalability compared to the traditional algorithms of DFS and BFS.
Q1Q2Q3Q4Q5Q6Q70050505050100100100100QueryExecution time (s) DFS + Automaton BFS + Automaton PathDB
Figure 4. Execution time of Queries over the SF1 Dataset.
Table 2. Execution time of Queries over the SF1 Dataset comparing the baseline with PathDB.
Query DFS + Automaton (s) BFS + Automaton (s) PathDB (s)
Q1 70.334 83.689 0.308
Q2 70.088 77.523 1.159
Q3 68.671 71.152 0.057
Q4 TimeOut TimeOut 2.439
Q5 67.423 69.859 0.039
Q6 69.280 70.534 0.436
Q7 68.362 73.062 0.035
4.2. Comparison with graph databases
For this comparison, we selected representative systems from commercial contexts. For that, we evaluated three graph database platforms: Kùzu (kuzu2023, ), a high-performance native graph engine; Neo4j (neo4jdocs, ), one of the most widely adopted systems in the industry; and Memgraph (memgraphdocs, ), a streaming-first in-memory graph database focused on low-latency execution and support for both transactional and analytical workloads. These systems were selected based on their support for path queries, and relevance in the current graph data ecosystem.
We generated a set of queries based on Regular Path Queries (RPQ). We created 30 types of abstract regular expressions (e.g., A.B for concatenation), and for each abstract regular expression, we generated between 2 and 6 RPQ’s (e.g., likes.hasCreator). For each query, we selected a source node with an outdegree situated at the median with respect to the adjacent label.
The distribution of the abstract regular expressions and the number of queries generated per expression can be seen in Table 3.
Tested systems
Since PathDB supports multiple path evaluation semantics and operates intrinsically in-memory, it was necessary to select various commercial systems for comparison. It is important to note that systems were selected that operate in-memory, on disk, or both, with the latter configured primarily for in-memory operation. The systems used include:
  • Kuzu Version 0.7.0 (Kuzu);
  • Memgraph Version 2.21.0 (MemGraph);
  • Neo4J Community Edition Version 5.26.0 (Neo4J).
For the aforementioned systems, it is worth noting that Kuzu and MemGraph support ALL WALKS, Kuzu and Neo4J support TRAIL, and Kuzu support ACYCLIC.
Table 3. Abstract Query Type distribution
Abstract
Query Type
Num.
Types
Example
A.B 6 hasModerator.knows
A.B.C 6 hasModerator.knows.isLocatedIn
A+.B 6 (replyOf+).hasTag
A.B+ 6 replyOf.(replyOf+)
[…]
(A.B)+ 2 (likes.hasCreator)+
[…]
A+ 3 knows+
A*.B 6 (knows*).likes
A.B* 6 hasMember.(knows*)
A* 3 knows*
(A.B)* 2 (likes.hasCreator)*
(A.B)? 6 (knows.likes)?
A.B? 6 hasMember.(knows?)
A?.B 6 (hasCreator?).isLocatedIn
Abstract
Query Type
Num.
Types
Example
[…]
A? 6 hasCreator?
(A?)? 6 (likes?)?
[…]
Total query types = 166
Test results
Each query was run three times on the SF1 dataset, with the number of returned results limited to 100 entries and a timeout threshold of 120 seconds per execution.
To evaluate system performance, we executed a set of 166 queries covering a variety of path patterns. These queries were grouped into 30 abstract query types, covering a diverse range of path queries. The results obtained from this evaluation provide insight into execution behavior across different graph engines, and form the basis for the comparative analysis presented in the following section.
A.BA.B.CA+.BA.B+A—B+(A.B)+A.(B—C)A+A*.BA.B*A*(A.B)*(A.B)?A.B?A?.BA—BB—A(A.B)—CA—(B.C)(A—B)—CA+—BA*—BA?—BA—B?A?(A?)?A—(B—C)(A—B)+(A—B)?(A—B)*001111222233334444Abstract Query PatternExecution time (s)PathDBKuzuMemGraph
Figure 5. Average Execution Time in Seconds by Query Type with Walk Semantic. Missing points indicate unsupported queries.
A.BA.B.CA+.BA.B+A—B+(A.B)+A.(B—C)A+A*.BA.B*A*(A.B)*(A.B)?A.B?A?.BA—BB—A(A.B)—CA—(B.C)(A—B)—CA+—BA*—BA?—BA—B?A?(A?)?A—(B—C)(A—B)+(A—B)?(A—B)*001111222233334444Abstract Query PatternExecution time (s)PathDBKuzuNeo4J
Figure 6. Average Execution Time in Seconds by Query Type with Trail Semantic. Missing points indicate unsupported queries.
A.BA.B.CA+.BA.B+A—B+(A.B)+A.(B—C)A+A*.BA.B*A*(A.B)*(A.B)?A.B?A?.BA—BB—A(A.B)—CA—(B.C)(A—B)—CA+—BA*—BA?—BA—B?A?(A?)?A—(B—C)(A—B)+(A—B)?(A—B)*001111222233334444Abstract Query PatternExecution time (s)PathDBKuzu
Figure 7. Average Execution Time in Seconds by Query Type with Simple Semantic. Missing points indicate unsupported queries.
The results of the experiments demonstrate that PathDB has a significant advantage over the other systems tested: it can execute all types of path queries, something that other systems such as Memgraph or Kuzu cannot do. This is clear in Figures 5, 6, and 7, which show the execution times for different types of graph searches.
Unlike other systems that only work well with certain specific queries, PathDB consistently handles all path query cases presented. For example, in Figure 5 we see how it responds well even to complex queries that cause other systems to fail. Of course, there is a small cost in speed. But PathDB offers a much better balance between speed and the ability to handle any type of search.
Most interestingly, PathDB does not require special modifications or extra implementations to work with difficult queries, something that other systems do require. Figures 6 and 7 clearly show this advantage, especially when working with searches that repeat patterns or alternate between different paths.
5. Conclusion and Future work
PathDB, while being an easy-to-use system with several configurable options and a query language tailored to paths, is still under development. Nevertheless, its architecture and applied optimizations have shown that it delivers acceptable performance compared to baseline algorithms, providing fast responses. Future work will focus on researching and implementing new optimization techniques, as well as exploring different storage methods to extend its capacity.
Acknowledgements.
This work was supported by ANID FONDECYT Chile through grant 1221727. R. García was supported by CONICYT-PFCHA / Doctorado Nacional / 2019-21192157.
Appendix A Query Times
Table 4. Execution Times (s) by Query Pattern Across All Semantics
Query Walk Trail Simple
PathDB Kuzu MemGraph PathDB Kuzu Neo4J PathDB Kuzu
A.B 0.609 0.253 0.407 0.659 0.179 0.002 0.892 0.179
A.B.C 0.902 0.341 0.286 0.944 0.267 0.002 1.055 0.255
A+.B 0.491 0.287 0.318 0.466 0.201 0.002 0.506 0.230
A.B+ 0.979 1.926 0.288 1.479 1.736 0.003 1.004 1.697
[…]
(A.B)+ 3.305 2.439 3.650
[…]
A+ 0.521 0.057 0.435 0.547 0.039 0.002 0.619 0.040
A*.B 1.130 0.277 0.319 0.939 0.370 0.002 1.233 0.324
A.B* 1.913 2.228 0.289 1.648 2.110 0.003 2.490 1.993
A* 1.062 0.072 0.439 0.872 0.051 0.002 1.201 0.049
(A.B)* 3.706 2.554 3.694
(A.B)? 0.852 0.764 0.829
A.B? 0.637 0.720 0.411 0.785 0.522 0.001 0.659 0.513
A?.B 0.694 0.157 0.409 0.695 0.088 0.001 0.875 0.097
[…]
A? 0.646 0.053 0.513 0.029 0.001 0.694 0.031
(A?)? 1.173 1.184 1.265
[…]
Table 5. Number of Answered and Unanswered queries per Database (regardless of semantics)
Database Answered Queries Unanswered Queries
PathDB 30 0
Kuzu 19 11
Neo4J 19 11
MemGraph 18 12

### v2

@@
 This modular design allows for targeted optimizations and modifications without impacting overall functionality.
-Benchmark experiments illustrate PathDB’s execution times and flexibility in handling dynamic and complex path queries, compared to baseline methods like Depth-First Search (DFS) and Breadth-First Search (BFS) guided by an automaton, highlighting its optimizations that contribute to its performance.
+Benchmark experiments illustrate PathDB’s execution times and flexibility in handling dynamic and complex path queries, compared to baseline methods like Depth-First Search (DFS) and Breadth-First Search (BFS) guided by an automaton, highlighting PathDB optimizations that contribute to its performance.
+PathDB was also evaluated against leading commercial graph systems, including Neo4j, Memgraph, and Kùzu.
+Benchmark experiments demonstrated PathDB’s competitive execution times and its ability to support a wide range of path query types.
 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
@@
 While these systems support path querying, they share a common characteristic: the reliance on specific algorithms to process recursive paths queries.
-However, Neo4j, Kuzu, and Memgraph offer only limited support for path queries.
-PathDB focuses on utilizing execution trees instead of recursive algorithms, as used by most systems.
-This approach allows for future optimization of these trees through techniques such as predicate pushdown and other enhancements.
+In contrast, Neo4j, Kuzu, and Memgraph offer only limited support for path queries.
+PathDB is a Java-based graph database designed for in-memory data loading and querying.
+By utilizing Regular Path Queries (RPQ) and a closed path algebra.
+Instead of relying on traditional graph traversal algorithms such as Breadth-First Search (BFS) or Depth-First Search (DFS), PathDB adopts a recursive query evaluation strategy centered around the use of the join operator.
+As a result, query evaluation in PathDB can be represented as execution trees, this approach allows for future optimization of these trees through techniques such as predicate pushdown and other enhancements.
 Its uniqueness lies in a path manipulation algebra, inspired by relational algebra, based on the concept of sets of paths.
@@
 It is worth noting that all these operators function exclusively on sets of paths.
-PathDB’s design features three main components: the parser, the logical plan, and the physical plan.
-By utilizing a path query pattern inspired by the GQL standard (2024iso, ), PathDB enables users to query paths and return a set of paths as results.
+PathDB’s design features three main components: the parser that perform the query validation and extraction, the logical plan responsible for converting queries into a logical plan and applying subsequent optimizations, and the physical plan which extracts the paths to be consulted using the different algebra operators.
+PathDB’s query language uses a path pattern inspired by the GQL standard (2024iso, ), allowing users to query paths and retrieve sets of paths.
 The primary objective of this paper is to showcase the innovative features and capabilities of PathDB.
-We provide a comprehensive overview of how graph data is stored, the query evaluation workflow, the syntax and operation of its query language, and related optimizations.
-Additionally, we demonstrate how PathDB’s design allows for targeted optimizations and modifications without disrupting overall functionality.
-This includes optimizations in the logical plan, changes to the physical plan, and the ability to seamlessly transition between different storage systems.
+We present a comprehensive overview of its graph data storage model, configuration parameters, and query interface.
+Furthermore, we describe the query evaluation workflow, detailing the syntax and operation of its query language as well as related optimizations.
+Additionally, we demonstrate how PathDB’s design allows for targeted optimizations and operator modifications without disrupting overall functionality.
 Section 2 provides a system overview, describing the architecture and data storage.
 Section 3 presents a demonstration of PathDB’s interface and usage, including a comparison with base methods like DFS and BFS with an automaton.
-Finally, Section 4 offers conclusions and future work.
+Finally, Section 4 presents the conclusions and the future work.
 Figure 1. A graph representing a social network (drawn from the LDBC SNB benchmark) (EDBT.2025.63, ).
@@
 In this section, we provide a comprehensive overview of the principal PathDB features.
-We will cover how the data is stored, the architecture’s design, the operation of its query language, and related optimizations.
+We cover how the data is stored, the architecture’s design, the query language features, and logical plan optimizations.
 It is worth noting that PathDB is a Java-based system designed to load and query data in memory.
@@
 By using a path query pattern based on Regular Path Queries (RPQ), PathDB enables users to query for paths and return a set of paths as results.
-Importantly, PathDB employs a closed path algebra, meaning that each operator takes a set of paths as input and returns a set of paths as output.
+Importantly, PathDB implements a closed path algebra, meaning that each operator takes a set of paths as input and returns a set of paths as output.
 Graph Storage
@@
 The information about nodes and edges is stored in a data structure that uses a compressed sparse row with vertical partitioning (CSR VP) as its base.
-Essentially, a CSR is represented by two lists: the first list maintains the offsets into the second array, marking the beginning of the outgoing edges for the corresponding vertex.
-The end of this list is marked by the beginning of the next vertex.
-The second array contains the edge identifiers (Tetzel2018, ).
-Combining CSR with vertical partitioning (VP) allows us to leverage the characteristics of a CSR while using labels, which is beneficial for our case since Regular Path Queries make use of these labels (Tetzel2018, ).
-In our implementation, we have two hashmaps: one where the key is the label and the value is a Node, and another where the key is the label and the value is a linked list of Edges that contain this label.
-It is worth mentioning that, so far, we have demonstrated how the graph is stored, but not how to work with paths.
-Since the logical and physical operators in PathDB handle sets of paths, it is necessary to transform the nodes and edges of a graph $G$ into sets of paths.
-To achieve this, two main functions are used: $\texttt{Paths}_{0}(G)$, which converts the set of nodes into paths of length 0, and $\texttt{Paths}_{1}(G,l)$, which transforms the edges of a graph $G$ with a specific label $l$ into a set of paths of length 1.
+Essentially, a CSR (Tetzel2018, ) is represented by two lists: the first list, often called offsets, stores the index positions that indicate where each vertex’s list of outgoing edges begins in the second list, as illustrated in Figure 4.
+The second array, commonly referred to as columns, contains the destination vertices of all edges, flattened into a single list.
+Combining CSR with vertical partitioning (VP) allows us to leverage the characteristics of a CSR while using labels, which is beneficial for our case since Regular Path Queries make use of these labels.
+It is important to note that when the graph contains too many labels, VP becomes inefficient: it creates many small tables, leading to more joins, more reads, and worse performance.
+Offsets0 2 5 6 Edge IDse0 e2 e3 e4 e5 e1 e6 e7
+Figure 4. Compressed sparse row CSR example.
+In our CSR based design, we define hashmaps: one where the key is the label and the value is a Node, and another where the key is an edge label and the value is a linked list of edges that contain this label.
+It is worth mentioning that, so far, we have shown how the graph is stored, but not how to work with paths.
+Since the logical and physical operators in PathDB handle sets of paths, we need to transform the nodes and edges of a graph $G$ into sets of paths.
+To achieve this, two main functions are used: $\texttt{Paths}_{0}(G)$, which retrieves the paths of length 0 of a graph $G$, and $\texttt{Paths}_{1}(G,l)$, which retrieve the paths of length 1 with a label $l$ from $G$.
 Path Algebra
-Our path algebra (EDBT.2025.63, ) consists of four fundamental operators, which can be composed and evaluated over graph-structured data.
+Our path algebra (EDBT.2025.63, ) defines five core operators that operate over sets of paths.
 The selection operator ($\sigma$) filters paths by applying conditions over nodes or edges.
 The union operator ($\cup$) merges two sets of paths, $a$ and $b$, into a single set.
-The path join operator ($\bowtie^{\tau}$) concatenates paths from set $a$ with compatible paths from set $b$; the join may optionally be constrained by a path semantic $\tau$, which can be one of isTrail for no repeated edges, isAcyclic for no repeated nodes, or isSimple for no repeated nodes, except the first and last node.
+The join operator ($\bowtie^{\tau}$) concatenates paths from set $a$ with compatible paths from set $b$; the join may optionally be constrained by a path semantic $\tau$, which can be one of isTrail for no repeated edges, isAcyclic for no repeated nodes, or isSimple for no repeated nodes, except the first and last node.
 Finally, the recursive operator ($\phi^{\tau}$) applies repeated join operations to compute transitive expansions of a path set, either until a fix point is reached or until specific criteria are satisfied.
 Like the join, the recursive operator can also be constrained by a path semantic $\tau$, allowing it to generate only trails, simple, or acyclic paths when required.
+Finally, the projection operator allows us to extract specific components from the resulting paths and optionally limit the number of results returned.
 These operators form the foundation for expressing rich navigational queries.
-An example that demonstrates the combined application of selection, union, join, and recursion is presented in Figure 1, shaped as an execution tree.
+An example that demonstrates the combined application of selection, union, join, recursion, and projection is presented in Figure 1, shaped as an execution tree.
 Path Query Pattern
-A path query pattern is an expression inspired by the GQL standard (2024iso, ), and consists of three main components: the MATCH clause, that defines the path pattern, the WHERE clause, that defines the selection conditions, and the RETURN clause, that specifies the output.
+A path query pattern is an expression inspired by the GQL standard (2024iso, ), and consists of three main components: the MATCH clause, that indicate the path pattern, the WHERE clause, that indicate the selection conditions, and the RETURN clause, that indicates the output terms.
         MATCH <restrictor>? <var> = <pathPattern>
@@
 Valid constructions for r include: a label $r_{1}$; a negated label $!r_{1}$; concatenation $r_{1}r_{2}$; alternation $r_{1}\mid r_{2}$; Kleene star $r_{1}^{*}$; positive closure $r_{1}^{+}$; and the optional expression $r_{1}?$.
-The WHERE clause consists of a set of selection condition terms used to filter the paths retrieved in the MATCH clause.
+The WHERE clause contains a selection condition used to filter the paths retrieved in the MATCH clause.
 Let $i\geq 1$ be a natural number, pr a property label, $v$ a value, $\oplus$ a comparison operator ($=$, $\neq$, $<$, $>$, $\leq$, $\geq$), and sVar and tVar symbolic variables representing the source and target nodes of a path pattern.
 A selection condition $c$ is defined recursively.
-A simple selection condition in PathDB can be any expression such as sVar.pr $\oplus$ v, tVar.pr $\oplus$ v, FIRST().pr $\oplus$ v, LAST().pr $\oplus$ v, NODE(i).pr $\oplus$ v, EDGE(i).pr $\oplus$ v, LABEL(sVar) $\oplus$ v, LABEL(tVar) $\oplus$ v, LABEL(FIRST()) $\oplus$ v, LABEL(LAST()) $\oplus$ v, LABEL(NODE(i)) $\oplus$ v, LABEL(EDGE(i)) $\oplus$ v, LENGTH() $\oplus$ v, as well as boolean predicates such as ISTRAIL(), ISSIMPLE(), and ISACYCLIC().
+A selection condition in PathDB can be any expression such as:
+  • sVar.pr $\oplus$ v
+  • tVar.pr $\oplus$ v
+  • FIRST().pr $\oplus$ v
+  • LAST().pr $\oplus$ v
+  • NODE(i).pr $\oplus$ v
+  • EDGE(i).pr $\oplus$ v
+  • LABEL(sVar) $\oplus$ v
+  • LABEL(tVar) $\oplus$ v
+  • LABEL(FIRST()) $\oplus$ v
+  • LABEL(LAST()) $\oplus$ v
+  • LABEL(NODE(i)) $\oplus$ v
+  • LABEL(EDGE(i)) $\oplus$ v
+  • LENGTH() $\oplus$ v
+Additionally boolean predicates such as ISTRAIL(), ISSIMPLE(), and ISACYCLIC() can be applied to filter path results.
 If $c_{1}$ and $c_{2}$ are selection conditions, then $(c_{1}{\small\texttt{ AND }}c_{2})$ and $(c_{1}{\small\texttt{ OR }}c_{2})$ are complex selection conditions.
-Finally, the RETURN clause specifies the elements of the path to be projected in the result.
+Finally, the RETURN clause is used to specify the elements of the path to be projected in the result.
 Projection expressions may reference any variable defined in the MATCH clause, including path and node variables.
@@
 It can also include functions that extract structural or semantic features of the path, such as LABEL(), LENGTH(), ISTRAIL(), ISSIMPLE(), and ISACYCLIC().
-An optional LIMIT clause may be appended to restrict the number of returned paths.
+An optional LIMIT clause may be appended to restrict the number of returned results.
 System Design
-The PathDB design is structured around three primary components, reflecting its query evaluation workflow, as shown in Figure 2: the parser, the logical plan, and the physical plan.
+The PathDB design is structured around three primary components, reflecting its query evaluation workflow, as shown in Figure 2: the parser, the logical planner, and the physical planner.
 First, the parser is responsible for validation, extraction, and transformation of a path query pattern into an abstract syntax tree (AST).
@@
 The AST is then transformed into a logical plan tree.
-Essentially, we convert the regular expression and endpoint filters (represented as a tree) into multiple algebraic operators.
-It is worth noting that our logical plan tree is executed in a bottom-up manner, from the leaves to the root.
-For example, to retrieve paths that start at a node named $Moe$ and traverse one or more repetitions of either a Knows or Likes edge followed by a HasCreator edge, a user can formulate the query:
+Essentially, we convert the path query pattern into multiple algebraic operators.
+It is worth noting that our logical plan tree is evaluated in a bottom-up manner, from the leaves to the root.
+For example, to retrieve paths that start at a with a property $name=Moe$ and traverse one or more repetitions of either a Knows or Likes edge followed by a HasCreator edge, a user can formulate the query:
 MATCH TRAIL p = (x)-[(Knows|Likes.HasCreator)+]->(y) WHERE x.name = "Moe" RETURN y.name LIMIT 10.
 The above query is based on the graph of Figure 1, the corresponding logical plan is presented in Figure 3(a).
-In this logical plan, the tree initially retrieves all edges from the graph $G$ labeled with Knows, Likes and HasCreator, subsequently transforming these edges into a set of paths.
+This logical plan specifies the operators required to evaluate the query: it first retrieves all edges from the graph $G$ labeled with Knows, Likes and HasCreator, and then transforms these edges into a set of paths.
 If we evaluate the expression from the left-hand side of the union, a recursive operator is applied over the set of paths labeled with Likes.
@@
 It is important to note that the path semantic isTrail is enforced on the resulting paths—both for the recursion and for the join—ensuring that no edge is repeated during evaluation.
-Once the recursion completes, we filter the resulting paths to retain only those whose source node has the property "Name = Moe".
-Finally, we return the names of the individuals in the last node of each path, limiting the output to 10 results.
+Above the union operator, the logical plan applies a selection that retains only paths whose source node has the property "Name = Moe".
+Finally, a projection operator extracts the names associated with the last node of each remaining path, limiting the output to the top 10 results.
 As can be observed, since we are working with a tree structure where each operation processes a set of paths, various optimizations can be applied during execution, which could improve execution times.
-These optimizations are discussed further later on.
-Finally, the output of the logical plan is used to retrieve results from the graph through the physical plan.
-To achieve this, each logical operator in the logical plan is translated into a corresponding physical operator.
+Finally, the physical plan implements the necessary algorithms to execute the operations specified by the logical plan, retrieving the corresponding results from the graph.
+To achieve this, each logical operator in the logical plan is translated into a corresponding physical operator algorithm.
 In this case, the physical plan operates on the graph stored in memory, processing each result on demand.
-The design of PathDB allows for the isolation of these three components, thereby enabling optimizations or modifications to be made to each one without causing significant issues.
+The design of PathDB allows for the isolation of these three components, thereby enabling optimizations or modifications to be made to each one without causing significant issues for the rest of the operator.
 This isolation, for instance, allows for optimizations in the logical plan, changes to the physical plan, such as moving from memory storage to disk storage or any other system, like a database, among other possible modifications.
@@
 The algebraic framework enables the systematic transformation of queries into more efficient forms while preserving their semantic integrity.
-A prominent optimization technique within this context is predicate pushdown, which involves moving constraints on the source or target nodes of paths closer to the leaf nodes in the logical plan.
+A prominent optimization technique within this context is predicate pushdown (10.5555/645920.672839, ), which involves moving constraints on the source or target nodes of paths closer to the leaf nodes in the logical plan.
+Predicate pushdown works well in practice because it minimizes the amount of data that flows through the query plan.
+By applying filtering conditions as early as possible.
 For instance, for the query discussed before, MATCH TRAIL p = (x)-[(Knows|Likes.HasCreator)+]->(y) WHERE x.name = "Moe" RETURN y.name LIMIT 10, Figure 3(a) illustrates the original query plan.
@@
 In this section, we demonstrate PathDB’s interface and usage.
-We showcase how to perform path queries, load data into memory, and interact with the system’s components.
+We showcase how to load data into memory, perform path queries, and interact with the system’s components.
 Additionally, we provide a comparison with base methods such as Depth-First Search (DFS) and Breadth-First Search (BFS) with an automaton to illustrate PathDB’s capabilities.
@@
 A.BA.B.CA+.BA.B+A—B+(A.B)+A.(B—C)A+A*.BA.B*A*(A.B)*(A.B)?A.B?A?.BA—BB—A(A.B)—CA—(B.C)(A—B)—CA+—BA*—BA?—BA—B?A?(A?)?A—(B—C)(A—B)+(A—B)?(A—B)*001111222233334444Abstract Query PatternExecution time (s)PathDBKuzuMemGraph
-Figure 5. Average Execution Time in Seconds by Query Type with Walk Semantic. Missing points indicate unsupported queries.
+Figure 6. Average Execution Time per Query Type with Walk Semantics. Missing bars indicate unsupported queries.
 A.BA.B.CA+.BA.B+A—B+(A.B)+A.(B—C)A+A*.BA.B*A*(A.B)*(A.B)?A.B?A?.BA—BB—A(A.B)—CA—(B.C)(A—B)—CA+—BA*—BA?—BA—B?A?(A?)?A—(B—C)(A—B)+(A—B)?(A—B)*001111222233334444Abstract Query PatternExecution time (s)PathDBKuzuNeo4J
-Figure 6. Average Execution Time in Seconds by Query Type with Trail Semantic. Missing points indicate unsupported queries.
+Figure 7. Average Execution Time per Query Type with Trail Semantics. Missing bars indicate unsupported queries.
 A.BA.B.CA+.BA.B+A—B+(A.B)+A.(B—C)A+A*.BA.B*A*(A.B)*(A.B)?A.B?A?.BA—BB—A(A.B)—CA—(B.C)(A—B)—CA+—BA*—BA?—BA—B?A?(A?)?A—(B—C)(A—B)+(A—B)?(A—B)*001111222233334444Abstract Query PatternExecution time (s)PathDBKuzu
-Figure 7. Average Execution Time in Seconds by Query Type with Simple Semantic. Missing points indicate unsupported queries.
+Figure 8. Average Execution Time per Query Type with Simple Semantics. Missing bars indicate unsupported queries.
 The results of the experiments demonstrate that PathDB has a significant advantage over the other systems tested: it can execute all types of path queries, something that other systems such as Memgraph or Kuzu cannot do.